A work cell in a local manufacturer has a workstation with 4 machines and each unit needs to pass through each machine in order to be completed. Each machine has a production capacity of four units per hour. What is the process time of the work cell in minutes per unit?
Answer #1

Process time= (1/capacity)*60

=(1/4)*60=15 min

So each machine has 4 min of process time.

Sp process cycle time= sum of process time=15+15+15+15

=60 min